Please help me solve the following question: If you total the ages of Bob and Bill, you will get 50. Ten years ago Bob was twice as old as Bill. How old are they now?
Here is what I have so far:
Bob (x)
Bill (y)
x+y=50
X-10=2(y-10)
Is this correct and how do I solve?

Answer by Alan3354(69443)

If you total the ages of Bob and Bill, you will get 50. Ten years ago Bob was twice as old as Bill. How old are they now?
Here is what I have so far:
Bob (x)
Bill (y)
x+y=50
X-10=2(y-10)
Is this correct and how do I solve?
------------------
Your equations are good. Getting the eqns right is usually the hard part.
---
x+y=50
X-10=2(y-10)
-----
x+y=50 --> y = 50-x
X-10=2(y-10) = 2y - 20
---
Sub for y in the 2nd eqn
x-10 = 2(50-x) - 20
x-10 = 100-2x - 20 = 80-2x
3x = 90
x = 30
y = 20